OpenStudy (anonymous):

Find the solution to the equation 64^(3 – x) = 4^2x. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

Let's start by stating that: \[64=4^{3}\] Now look at the problem: \[64^{3-x}=4^{2x}\] Substitute in the first statement: \[4^{3(3-x)}=4^{2x}\] \[4^{9-3x}=4^{2x}\] Now we take the log base 4 on each side to remove the 4. \[\log_{4} 4^{9-3x}=\log_{4} 4^{2x}\] Taking the log base 4 cancels the 4 on each side so we get: \[9-3x=2x\] Now we simplify: \[9=5x\] \[x=\frac{9}{5}\]